# Anisotropy of Nanohertz Gravitational Wave Background and Individual Sources from Supermassive Binary Black Holes Based on Cosmological Simulation

### v2

@@ 3.3.3 GW Sources in a Shell
We are interested in whether the distribution of SMBBHs can reproduce the large-scale structure of the universe, and thus be used for cosmological purpose like galaxies. Given that the GW sources are situated at different distances from us, it is difficult to determine the physical scale corresponding to specific angular scales. To better study the clustering patterns exhibited in the spatial distribution of the GW sources from angular power spectrum, we selected GW sources located in different spherical thin shells with radius from $R-dR$ to $R+dR$, where we choose $dR=10$ Mpc and radius $R$=382, 764 and 1528 Mpc, which corresponds to redshift $z=0.09$, $0.19$ and $0.41$ respectively. Figure 12 shows the angular position distribution sky map of these GW sources in the three different spherical shells (from top to bottom). As the radius of the spherical shell increases, more and more SMBBHs gather within the range of the shell, thus appearing increasingly crowded. In Figure 13 we also plotted a small patch of the celestial sphere, where we simultaneously displayed the distribution of SMBBHs and background galaxies. The distribution of SMBBHs is sparser than that of galaxies, particularly in spherical shells with smaller radius. It appears that SMBBHs alone may only provide incomplete information about the large-scale structure of the universe.
@@
In order to make a quantitative estimation about to what extent could the distribution of SMBBHs trace the underling matter distribution of the universe, we calculated the $C_{l}$ coefficients for the number density of both GW sources and galaxies in these spherical shells and show the results in Figure 14. Different $l$ corresponds to different angular scale $\alpha$ based on the relationship $\alpha=\frac{\pi}{l}$. When the radius $R$ of the spherical shells is fixed, a certain angular scale $\alpha$ corresponds to a specific physical scale $R\alpha$. We observe that, compared to SMBBHs, the $C_{l}/C_{0}$ values of galaxy distribution exhibit more significant fluctuations when $l$ is relatively small, which indicates the stronger clustering of galaxies on large scales. In contrast, $C_{l}/C_{0}$ of SMBBH distribution show less variation within the corresponding range. On the other hand, the behavior of $C_{l}/C_{0}$ of both galaxies and SMBBHs gradually converges as $l$ increases. This indicates that only on relatively smaller scales, the distribution of SMBBHs may exhibit similar patterns to that of galaxies.
@@
We further calculated the angular two point correlation function $\omega(\alpha)$ in these three shells as defined by the following formula:
@@
where $DD(\alpha)$ and $RR(\alpha)$ correspond to the number of SMBBH or galaxy pairs with angular separation $\alpha$ in data-data and random-random catalogs, respectively, whereas $DR(\alpha)$ corresponds to the number of pairs with separation $\alpha$ calculated between data and random catalogs. The results are shown in Figure 15. In the figures, we observe that the angular correlation functions exhibit similar trends to those of the angular power spectra. For smaller spherical shells and smaller angular ranges, the correlation functions of SMBBHs and galaxies are relatively close to each other. However, for larger spherical shells or larger angles, the correlation functions of the two show significant deviations, where the correlation of galaxies is stronger than that of SMBBHs on the corresponding angular scales. This is understandable, because as can be clearly seen from Figure 12 and Figure 13, the distribution of SMBBHs is more homogeneous than that of galaxies. Therefore, they can only reproduce the clustering pattern of galaxies on small scales, while the more densely populated regions of galaxies on large scales are not reflected in the distribution of SMBBHs.
Of course, these are ideal theoretical discussions within the simulated universe. In reality, we need to be able to distinguish these SMBBHs and their angular positions on the celestial sphere. With the improvement of experimental techniques in the future, we will be able to detect an increasing number of individual sources, by combining them with the distribution of stellar-mass black hole events, and calculate their correlation functions or cross-correlations functions with galaxies, the clustering properties of binary black hole events still hosts the possibility to provide us with useful information across larger scale ranges.
@@
Figure 12: The angular distribution sky map of GW sources in spherical shells with radius $R=$ 382, 764 and 1528 Mpc, respectively (from top to bottom), where red point represents the position of each GW source.
@@
Figure 13: A small portion of the sky map with both GW sources and galaxies in spherical shells with radius $R=$1528, 764 and 382 Mpc, respectively (from left to right). The red points represent the positions of SMBBHs, while the green points represent the positions of galaxies.
@@
Figure 14: Angular power spectrum coefficients $C_{l}$ as a function of $l$ for the number density of GW sources (top) and galaxies (bottom) in different spherical shells.
@@
Figure 15: The two-point correlation functions $\omega$ of galaxies and SMBBHs as a function of $\alpha(^{\circ})$ for spherical shells with radius $R=382$, $764$ and $1528$ respectively (from left to right).
4 Detectabilities
@@
Table 1: Parameter settings for two PTAs. Columns from left to right denote the notation of the PTA, total number of MSPs monitored by the PTA $N_{\rm p}$, the timing precision $\sigma_{\rm t}$, the total observational period $T_{\rm obs}$ and the mean cadence $\Delta t$ (Guo et al., 2022; Chen et al., 2023; Guo et al., 2024).
@@
4.1 Individual Sources
In the previous analysis, we assumed that the superposition of GW signals emitted by all SMBBH systems within the simulated light cone together form the GWB signal, which is correct for current PTAs. Actually, some individual GW sources can be detected by more powerful PTAs with enough SNR $\varrho$ ($\varrho>8$) in the future. So these individual GW sources can be resolvable and separated from background. For future CPTA and SKA-PTA featuring greater detection capability and higher accuracy (The parameters related to their experimental configurations are presented in Table 1), assuming a 20-years’ observation period, we calculated the SNR of our sampling SMBBHs based on Equation (85) in Guo et al. (2022), and selected those GW sources that can be resolved by these future PTAs. We show the three dimensional spatial distribution of these selected individual sources in Figure 16 for both CPTA and SKA-PTA, where we also use colors to represent the SNR of each source. Our results suggest that, future CPTA is projected to detect 1286 individual GW sources with SNR $\varrho>8$ during 20 years’ observation time, while SKA-PTA is expected to identify 22695 individual GW sources with the same observation duration (See also Chen et al. 2023. The ranges predicted by this paper cover our results).
The statistical distribution of SNRs of all individual GW sources for CPTA (red) and SKA-PTA (green) are shown in Figure 17. Their SNR distribution basically follow power law distribution. From the figure, one can observe the highest achievable SNR for the most optimistic individual cases. CPTA is capable of detecting individual sources with an SNR of over 100. Meanwhile, SKA-PTA, with its advanced sensitivity, is expected to detect GW sources with SNR values reaching more than 700.
Based on these resolvable individual GW sources by future PTAs, we can again use spherical harmonic expansion to analyze the angular power spectrum relating to their number density distributions across the celestial sphere. Since the number of detectable individual GW sources for CPTA is rather limited, we mainly focus on the individual GW sources that are detectable by SKA-PTA. We further plotted their two dimensional angular distribution sky map in Figure 19, and the result, i.e., the $C_{l}$’s as a function of $l$, are shown in Figure 21.
@@ 4.1 Individual Sources
Figure 16: The three dimensional spatial distribution of GW sources detectable by CPTA (left) and SKA-PTA (right), where colors represent the logarithmic SNR, i.e., $\log_{10}(\varrho)$, of each GW source.
Figure 17: The logarithmic SNR distribution of individual GW sources for CPTA (red) and SKA-PTA (green). The distribution of SNR after considering the lensing effect is also shown as light red and light green lines.
Figure 18: The angular position distribution sky map of individual GW sources detectable by CPTA, where colors represent the comoving distance to us.
Figure 19: The angular position distribution sky map of individual GW sources detectable by SKA-PTA, where colors represent the comoving distance to us.
Figure 20: Angular power spectrum coefficients $C_{l}$ as a function of $l$ for the number density of the GW sources detectable by CPTA, where error bar indicate the cosmic variance.
Figure 21: Angular power spectrum coefficients $C_{l}$ as a function of $l$ for the number density of the GW sources detectable by SKA-PTA, where error bar indicate the cosmic variance.
4.2 Angular Resolution
In order to probe the large scale structure through GW signal from SMBBHs, we could have two distinct methodological options: one approach is to detect the GWB signal and study its anisotropy, while the other one envolves the detection of a multitude of individual sources with enough SNR and high precision locations, so that sky map of these individual sources can be constructed to study the underlying cosmic structures. As discussed in Guo et al. (2022) (see also Lee et al. 2011), it is possible that, under ideal conditions and with the incorporation of pulsar terms, PTAs utilizing a substantial ensemble of MSPs may achieve a localization precision as fine as $1^{\circ}$, thus the second approach for the detection of large-scale structures is anticipated to be within reach in the future.
@@ 5 Lensing Effect
The gravitational wave signals from distant SMBBHs may be affected by the lensing effects of foreground galaxies or dark matter halos as they propagate towards Earth. In this section, we will discuss whether the gravitational lensing effect will have a significant impact on the various results related to GWB and individual sources that we have derived in previous sections. We adopt the lensing model used in Oguri (2018), where they developed a hybrid framework to compute the magnification probability distribution function (PDF) that takes into account both weak and strong gravitational lensing effect for wide ranges of source redshifts and magnifications. We randomly assigned a magnification factor $\sqrt{\mu}$ to each gravitational wave event amplitude $h_{i}$ according to the PDF under this model. We compare the results with and without lensing effect in the following subsection.
5.1 Results
@@
Figure 22: A comparison of the isotropic and anisotropic properties of the GWB with and without considering gravitational lensing effects. Left panel: comparison of the GWB characteristic amplitude and spectrum (upper panel) and their difference (lower panel), lensing effects induced variation is one to three order of magnitude smaller than the original value; Right panel: comparison of the $C_{l}$ cofficients (upper panel) and their differences (lower panel), lensing effects induced variation is one order of magnitude smaller than the original value.
@@
From Figure 22 we can see that, for the characteristic amplitude of the GWB, upon incorporating the lensing effect into our calculations, we observed a subtle change in the characteristic amplitude of GWB. Specifically, the amplitude exhibited a slight increase at the order of $10^{-17}$, which is one to three order of magnitude smaller than the value of the amplitude itself. This increase is sufficiently small that it does not significantly alter the overall behavior of the GWB signal. For the anisotropy property of the GWB signal, lensing effects induced variation of the $C_{l}$ coefficients is roughly within two percent of the original value, the effect of gravitational lensing is not significant either.
For the detectability of individual sources, after incorporating the lensing effect, the number of resolvable individual sources for SKA increased from 22,695 to 22,828, and for CPTA, it increased from 1,286 to 1,287. This indicates an overall increasing trend, although the increase accounts for only a small fraction. The distribution of the SNR also shows only a subtle impact due to the lensing effect. This suggests that while gravitational lensing does have an influence on the detection of gravitational wave sources, its effect on the overall observability and SNR distribution is rather limited. In Figure 17, we present a comparison of the SNR distributions with and without the lensing effect, where lighter colors illustrate the distribution of SNR after considering the lensing effect.
@@ 6 Conclusions and Discussion
Utilizing data from numerical cosmological simulation, we generated more than $3\times 10^{7}$ SMBBHs distributed in the observable universe as GW sources of PTA. While maintaining the features of cosmic large-scale structure in the mock universe, we studied the isotropic and anisotropic characteristics of SMBBH sources and GWB signal, predicted the properties of detectable individual GW sources by CPTA and SKA-PTA, and also examined the influence of gravitational lensing effects.
@@
For GWB, we calculated its characteristic amplitude and spectrum, and found that the resulting GWB signal has an amplitude $A\approx 4.85\times 10^{-16}$ at the characteristic frequency ${\rm yr}^{-1}$. We demonstrated the anisotropy properties in the distribution of simulated GW sources and the GWB signal using sky maps, and derived the angular power spectrum coefficients $C_{l}$ as a function of $l$ for the energy density of GWB. we predict that $C_{1}/C_{0}\approx 2.50\times 10^{-3}\pm 2.04\times 10^{-3}$ and $C_{2}/C_{0}\approx 2.48\times 10^{-3}\pm 1.57\times 10^{-3}$. For angular power spectrum in different frequency bins, $C_{1}/C_{0}$ at higher frequency end can even reach $\sim 0.8$, while in the lower frequency end, $C_{l}/C_{0}$ generally ranges from $0.001$ to $0.01$ for $l>0$. We investigated the clustering patterns of GW sources in spherical shells at various redshifts and discovered that, on relatively smaller scales, the clustering patterns of SMBBHs are more similar to that of galaxies. If integrated with data from stellar-mass black hole events, this approach holds the potential to more comprehensively reflect the cosmic large-scale structure. We also studied the detection capability of the future CPTA and SKA-PTA for individual sources. From our calculation results, we predict that CPTA is projected to detect 1286 individual GW sources with a SNR greater than 8 during 20 years’ observation time, while SKA-PTA is projected to detect as many as 22695 individual GW sources, some of which may even achieve a SNR exceeding 700. These mock individual sources can provide more realistic samples to further study. Finally, we took gravitational lensing effects into account and found that lensing effects slightly enhanced our results, including the characteristic amplitude of the GWB signal and the number of detectable individual sources. However, overall, the impact is rather limited.
@@
Below, we offer some further discussions. As was mentioned at the end of Section 3.2, our result of the GWB characteristic spectrum lies below the current experimental data points. This discrepancy may stem from several factors. One such factor is that SAM inevitable has its own uncertainties in its model construction. If we consider other versions of SAM that have slightly higher merger rates, e.g., Henriques2015 (Yang et al., 2019), it is still possible that SAM may offer results that is nearly compatible with observational data points. Furthermore, it is also possible that the GWB signal detected is a complex superposition of signals from multiple sources across the universe. While SMBBHs are known to be significant contributors to this background, they are not the sole constituents. Other astrophysical phenomena, such as binary white dwarf or the early universe’s phase transition processes, may also imprint their signatures on the GWB, thereby complicating the interpretation of the observed signals and may also contribute to the discrepancy. However, it is important to emphasize that, despite these complexities, our current projected value of the GWB amplitude falls well within the theoretical range of [$1\times 10^{-16}$, $5\times 10^{-15}$] given by existing literature.
Since current $N$-body simulations combined with semi-analytical models cannot provide accurate information about the frequencies and orbits of SMBBH systems, we also made the simplifying assumption that the orbits of the SMBBHs are perfectly circular when calculating the frequency distribution probability and GW strains. If future numerical simulations can provide more detailed information on the orbital dynamics and evolution of SMBBHs, we will be able to develop more precise models for PTA band signal prediction. We will also update our results in future works based on certain orbital eccentricity assumptions.
We have theoretically explored the possibility of reconstructing the large-scale structure of the universe using supermassive black holes. But it is important to recognize that the individual sources detected by PTAs may be subject to observational biases. These sources are likely to be found in large elliptical galaxies, which could introduce bias compared to the general galaxy population (Chen et al., 2020). To effectively utilize these individual sources for investigating the large-scale structure of the universe, it is crucial to gain a clear understanding of this bias and to implement appropriate corrections within our observational methodologies.

### v3

@@ 3.3.2 GW Sources in Shells
We are interested in whether the distribution of SMBBHs can reproduce the large-scale structure of the universe, and thus be used for cosmological purpose like galaxies. Given that the GW sources are situated at different distances from us, it is difficult to determine the physical scale corresponding to specific angular scales. To better study the clustering patterns exhibited in the spatial distribution of the GW sources from angular power spectrum, we chose one realization from the most optimal case with $\tau_{H}=0.1\,\mathrm{Gyr}$ and selected GW sources located in different spherical thin shells with radius from $R-dR$ to $R+dR$, where we choose $dR=10$ Mpc and radius $R$=382, 764 and 1528 Mpc, which corresponds to redshift $z=0.09$, $0.19$ and $0.41$ respectively.
Figure 10 shows the angular position distribution sky map of these GW sources in the three different spherical shells (from top to bottom). As the radius of the spherical shell increases, more and more SMBBHs gather within the range of the shell, thus appearing increasingly crowded. In Figure 11 we also plotted a small patch of the celestial sphere, where we simultaneously displayed the distribution of SMBBHs and background galaxies. The distribution of SMBBHs is sparser than that of galaxies. It appears that SMBBHs alone may only provide incomplete information about the large-scale structure of the universe.
@@
In order to make a quantitative estimation about to what extent could the distribution of SMBBHs trace the underling matter distribution of the universe, we calculated the $C_{l}$ coefficients with Healpix for the number density of both GW sources and galaxies in these spherical shells and show the results in Figure 12. Different $l$ corresponds to different angular scale $\alpha$ based on the relationship $\alpha=\frac{\pi}{l}$. When the radius $R$ of the spherical shells is fixed, a certain angular scale $\alpha$ corresponds to a specific physical scale $R\alpha$. We observe that, compared to SMBBHs, the $C_{l}/C_{0}$ values of galaxy distribution exhibit more significant fluctuations when $l$ is relatively small, which indicates the stronger clustering of galaxies on large scales. In contrast, $C_{l}/C_{0}$ of SMBBH distribution show less variation within the corresponding range. On the other hand, the behavior of the $C_{l}/C_{0}$ coefficients of SMBBHs and galaxies show more agreement at higher multipole moments (large $l$). The values of $C_{l}/C_{0}$ of SMBBHs in this range are slightly larger than those of galaxies, indicating marginally stronger clustering of SMBBHs on smaller scales. Moreover, the discrepancy between the $C_{l}/C_{0}$ values decreases with smaller shell radius, while they become nearly identical within the smallest shell. This indicates that only on relatively smaller physical scales, the distribution of SMBBHs may exhibit similar patterns to that of galaxies. Our results are basically consistent with the behavior of cross correlations between GWB and galaxies shown in Semenzato et al. (2024), where larger $l$ (smaller scale) leads to stronger correlations between the two and higher sigma-level detectability.
@@
We further utilized Corrfunc to calculate the angular two point correlation function $\omega(\alpha)$ in these three shells as defined by the following formula:
@@
where $DD(\alpha)$ and $RR(\alpha)$ correspond to the number of SMBBH or galaxy pairs with angular separation $\alpha$ in data-data and random-random catalogs, respectively, whereas $DR(\alpha)$ corresponds to the number of pairs with separation $\alpha$ calculated between data and random catalogs. The results are shown in Figure 13. In the figure, we observe that the angular correlation functions exhibit similar trends to those of the angular power spectra. For smallest spherical shells, the correlation functions of SMBBHs and galaxies are nearly identical to each other. However, for larger spherical shells or larger angles, the correlation functions of the two show more deviations. The angular correlation of SMBBHs is slightly stronger than that of galaxies on smaller angular scales, while the angular correlation of galaxies is stronger than that of SMBBHs on larger angular scales. The deviation behavior on large scales is understandable, because as can be clearly seen from Figure 10 and Figure 11, the distribution of SMBBHs is more homogeneous than that of galaxies. Therefore, they can only reproduce the clustering pattern of galaxies on small scales, while the more densely populated regions of galaxies on large scales are not reflected in the distribution of SMBBHs.
Of course, these are ideal theoretical discussions within the simulated universe, but these results may offer insights for exploring the correlations between the anisotropy of the GWB signal and the distribution of galaxies. We will offer more discussion on this point in the discussion section.
@@
Figure 10: The angular distribution sky map of GW sources in spherical shells with radius $R=$ 382, 764 and 1528 Mpc, respectively (from top to bottom) from one representative realization with $\tau_{H}=0.1\,\mathrm{Gyr}$, where red point represents the position of each GW source.
@@
Figure 11: A small portion ($\phi\in[0.7,1.7]$ rad, $\theta\in[0.2,0.9]$ rad) of the sky map with both GW sources and galaxies in spherical shells with radius $R=$382, 764 and 1528 Mpc, respectively (from left to right). The red points represent the positions of SMBBHs, while the green points represent the positions of galaxies.
@@
Figure 12: Angular power spectrum coefficients $C_{l}/C_{0}$ as a function of $l$ for the number density of GW sources (red lines) and galaxies (green lines) in different spherical shells with radius $R=382$, $764$ and $1528$ Mpc respectively (from up to down) from one representative realization with $\tau_{H}=0.1\,\mathrm{Gyr}$.
@@
Figure 13: The two-point correlation function $\omega$ of galaxies and SMBBHs as a function of angle $\alpha(^{\circ})$ for spherical shells with radius $R=382$, $764$ and $1528$ Mpc respectively (from left to right) from one representative realization with $\tau_{H}=0.1$ Gyr.
3.4 Individual Source Detectabilities
@@
Table 1: Parameter settings for several PTAs. Columns from left to right denote the notation of the PTA, total number of pulsars $N_{\rm p}$ observed by the PTA, the timing precision $\sigma_{\rm t}$, the total observational period $T_{\rm obs}$ and the mean cadence $\Delta t$ (Guo et al., 2022; Chen et al., 2023; Guo et al., 2025).
@@
In the previous sections, we subtracted resolvable bright individual sources for current IPTA when analysing the characteristics of GWB signal in order to better reflect the current observation status. It is expected that more individual GW sources can be detected by more powerful PTAs with enough SNR $\varrho$ ($\varrho>8$) in the future. So these individual GW sources can actually be resolvable and separated from background. For future CPTA and SKA-PTA featuring greater detection capability and higher accuracy (The parameters expected to achieve for their experimental configurations are presented in Table 1), assuming a 20-years’ observation period and considering unresolvable GWB as noise, we calculated the SNR of our sampling SMBBHs based on the following equation (Guo et al., 2022, 2025):
$\varrho^{2}\approx N_{\rm p}\frac{\chi^{2}h_{\rm c}^{2}}{h_{\rm n}^{2}},$ (8)
where
$h_{\rm n}(f)=\sqrt{fS_{\rm n,s}+h_{\rm b}^{2}},$ (9)
and $S_{\rm n,s}(f)=8\pi^{2}\sigma_{\rm t}^{2}f^{2}\Delta t$ is the power spectral density contributed by shot noise, and $h_{\rm b}$ is the unresolvable GWB signal we produced.
Our results show that, future SKA is projected to resolve a median of 76, 32, and 2 individual GW sources with SNR $\varrho>8$ during 20 years’ observation time for $\tau_{H}=0.1$, $5$, $10\,\mathrm{Gyr}$ repectively (See also Chen et al. 2023; Truant et al. 2025), while CPTA is expected to identify a median of 5, 3, and 0 individual GW sources with the same observation duration, substantially lower than SKA. We have also confirmed that no realization exceeds the upper bound for the number of resolvable sources in each frequency bin that was proposed in Boyle & Pen (2012). The distribution of the number of realizations versus the detectable individual GW source number counts by CPTA and SKA-PTA for all the three hardening timescales are shown in Figure 14, and the results for the median values and variance of the number of individual sources are summarised in Table 2. We also presented the three dimensional spatial distributions and two dimensional angular distribution of one representative realization for the most optimistic case with $\tau_{H}=0.1\,\mathrm{Gyr}$ in Figure 15 for SKA-PTA, where we use colors to represent the SNR and comoving distance of each source respectively. The statistical distribution of SNRs/frequencies of all resolvable individual GW sources by CPTA and SKA-PTA from three representative realizations with abundant, moderate, and sparse source counts are shown in Figure 16/Figure 17. From the figures, one can observe the highest achievable SNR in different cases. SKA-PTA, with its advanced sensitivity, is expected to detect GW sources with SNR values reaching around 100, whereas CPTA achieves notably lower SNRs, rarely exceeding 30. At the same time, SKA detects sources at frequencies greater than $10^{-8}$z in all scenarios, while CPTA typically detects sources within the $10^{-9}$z range.
Figure 14: The number of realizations versus the number count of GW sources detectable (SNR$>$8) by CPTA (red) and SKA-PTA (green) for $\tau_{H}=0.1$, $5$, $10\,\mathrm{Gyr}$ respectively (from left to right).
[…]
SNR$>$8
$\tau_{H}$/Gyr CPTA SKA-PTA
0.1 $5_{-4}^{+5}$ $76_{-38}^{+61}$
5 $3_{-2}^{+4}$ $32_{-16}^{+21}$
10 $0_{-0}^{+1}$ $2_{-2}^{+3}$
@@
Figure 15: The three dimensional spatial distribution (left) and the angular position distribution sky map (right) of individual GW sources detectable by SKA-PTA from one representative realization with $\tau_{H}=0.1\,\mathrm{Gyr}$, where colors represent the SNR and comoving distance of each source respectively.
Figure 16: The SNR distribution of resolvable individual GW sources for CPTA (upper panels) and SKA-PTA (lower panels) for $\tau_{H}=0.1,5,10\,\mathrm{Gyr}$ respectively (from left to right). Each panel presents results from three representative realizations with abundant, moderate, and sparse source counts, where $N_{\rm tot}$ denotes the total number of individual sources in each realization. The vertical black dashed lines indicate the SNR threshold at $\rho=8$.
Figure 17: The frequency distribution of resolvable individual GW sources for CPTA (upper panels) and SKA-PTA (lower panels) for $\tau_{H}=0.1,5,10\,\mathrm{Gyr}$ respectively (from left to right). Each panel presents results from three representative realizations with abundant, moderate, and sparse source counts, where $N_{\rm tot}$ denotes the total number of individual sources in each realization. The vertical black dashed lines indicate the lower frequency limit when total observation duration is 20 years.
@@ 4 Lensing Effect
The gravitational wave signals from distant SMBBHs may be affected by the lensing effects of foreground galaxies or dark matter halos as they propagate towards Earth. In this section, we will discuss whether the gravitational lensing effect will have a significant impact on the various results related to GWB and individual sources that we have derived in previous sections. We adopt the lensing model used in Oguri (2018), where they developed a hybrid framework to compute the magnification probability distribution function (PDF) that takes into account both weak and strong gravitational lensing effect for wide ranges of source redshifts and magnifications. We randomly assigned a magnification factor $\sqrt{\mu}$ to each gravitational wave event amplitude $h_{i}$ according to the PDF under this model. We recalculated the GWB isotropic and anisotropic properties with lensing effect, and compared the derived results with previous results in Figure 18 and Figure 19.
@@
Figure 18: A comparison of the GWB characteristic amplitude with and without considering gravitational lensing effects from one representative realization for $\tau_{H}=0.1\,\mathrm{Gyr}$ (left), $5\,\mathrm{Gyr}$ (middle) and $10\,\mathrm{Gyr}$ (right) respectively. The lower panels show the differences in the characteristic amplitude between the two cases. Lensing effects induced variation is generally two to three orders of magnitude smaller than the original value.
Figure 19: A comparison of the angular power spectrum of the GWB total energy density with and without considering gravitational lensing effects from one representative realization for $\tau_{H}=0.1$ Gyr (left), $5$ Gyr (middle) and $10$ Gyr (right) respectively. The lower panels show the differences in the $C_{l}/C_{0}$ coefficients between the two cases. Lensing effects induced variation is generally two to three orders of magnitude smaller than the original value.
@@
From Figure 18 we can see that, for the characteristic amplitude of the GWB, upon incorporating the lensing effect into our calculations, we observed a subtle change in the characteristic amplitude of GWB. Specifically, the amplitude exhibited a slight increase at the order of $10^{-17}$ or $10^{-18}$, which is two to three orders of magnitude smaller than the value of the amplitude itself. This increase is sufficiently small that it does not significantly alter the overall behavior of the GWB signal. For the anisotropy property of the GWB signal, lensing effects induced variation of the angular power spectrum of total energy density is shown Figure 19. We can see that, lensing induced variation on the $C_{l}$ coefficients is roughly at the order of $10^{-4}$, which is also two to three orders of magnitude smaller than the original value, the effect of gravitational lensing is not significant either. For the detectability of individual sources, the lensing effect can slightly increase the number of individual sources with SNR above 8 in some realizations. However, since the main limitation on the number of resolvable individual sources is the upper limit in each frequency bins, the overall impact of lensing on the individual source number count is negligible across all cases.
@@ 5 Conclusions and Discussion
Utilizing data from numerical cosmological simulation, we generated multiple realizations of SMBBHs distributed in the observable universe as GW sources of PTA. While maintaining the features of cosmic large-scale structure in the mock universe, we studied the isotropic and anisotropic characteristics of SMBBH sources and GWB signal accounting for cosmic variance through different realizations, predicted the properties of resolvable individual GW sources by CPTA and SKA-PTA, and also examined the influence of gravitational lensing effects.
@@
For GWB, we calculated its characteristic amplitude and spectrum, and found that the characteristic amplitude $h_{c}$ at frequency $f=1~{}\mathrm{yr}^{-1}$ is $4.08\times 10^{-16}$, $3.38\times 10^{-16}$, and $2.19\times 10^{-18}$ for $\tau_{H}=0.1\,\mathrm{Gyr}$, $5\,\mathrm{Gyr}$, and $10\,\mathrm{Gyr}$, respectively; while at $f=0.1\,\mathrm{yr}^{-1}$, $h_{c}=1.96\times 10^{-15}$, $1.42\times 10^{-15}$, and $3.13\times 10^{-16}$ for the same $\tau_{H}$ values in corresponding order. We demonstrated the anisotropy properties in the distribution of simulated GW sources and the GWB signal using sky maps, and derived the angular power spectrum coefficients $C_{l}$ as a function of $l$ for both the total energy density of GWB and energy density in different frequency bins. In the total energy case, our results show that the directional dependence of frequency distribution represents a more dominant source of anisotropy compared to the contribution from bright individual sources. We predict that for most of the realizations in the cases of hardening timescales of $0.1,5,$ and $10\,\mathrm{Gyr}$, the distribution of $C_{1}/C_{0}$ falls within the range of $[3.0\times 10^{-2},4.3\times 10^{-1}]$, $[5.3\times 10^{-2},6.8\times 10^{-1}]$, $[7.7\times 10^{-2},6.2\times 10^{-1}]$ respectively (16%-84% distribution range), and are insensitive to the substraction of additional bright sources. Regarding the angular power spectrum of energy density in different frequency bins, $C_{1}/C_{0}$ shows great fluctuations at higher frequencies, while the lowest-frequency bin demonstrates a lower level of anisotropy compared to the total energy density case, and the differences in the $C_{l}/C_{0}$ distributions under the three hardening timescales become more pronounced. In the lowest frequency bin, our results show that $C_{1}/C_{0}$ falls mostly within the range of $[2.2\times 10^{-3},3.5\times 10^{-2}$], $[4.0\times 10^{-3},6.0\times 10^{-2}]$, and $[2.1\times 10^{-2},2.9\times 10^{-1}]$ for $\tau_{H}=0.1$, $5$, and $10\,\mathrm{Gyr}$ respectively (16%-84% distribution range), and the substraction of additional bright sources will further lower the level of anisotropy in the $\tau_{H}=0.1$ and $5\,\mathrm{Gyr}$ cases to $[1.2\times 10^{-3},9.1\times 10^{-3}$] and $[2.7\times 10^{-3},2.0\times 10^{-2}]$. Furthermore, We investigated the clustering properties of GW sources in spherical shells at various redshifts and discovered that, on relatively smaller scales, the clustering patterns of SMBBHs are more similar to that of galaxies. Since the GWB should be predominantly contributed by relatively nearby sources, this distribution pattern of GW sources may also be reflected in the GWB signal characteristics. We also studied the detection capability of the future CPTA and SKA-PTA for individual sources. From our calculation results, SKA-PTA is expected to identify a median of 76, 32, and 2 individual GW sources with SNR $\varrho>8$ during 20 years’ observation time for $\tau_{H}=0.1$, $5$, and $10\,\mathrm{Gyr}$ respectively, while future CPTA is projected to resolve a median of 5, 3, and 0 individual GW sources with the same observation duration. These mock individual sources can provide more realistic samples for further studies. Finally, we took gravitational lensing effects into account and found that lensing effects on the characteristics of the GWB signal and the number of detectable individual sources are rather limited.
@@
Below, we offer some further discussions. As was mentioned at the end of Section 3.2, the average value of the GWB characteristic amplitude from our work, even in the most optimistic case with $\tau_{H}=0.1$ Gyr, lies below the current experimental data points. This discrepancy may stem from several factors. One such factor is that SAM inevitable has its own uncertainties in its model construction. If we consider other versions of SAM that have slightly higher merger rates, e.g., Henriques 2015 or Izquierdo-Villalba 2019 (Yang et al., 2019; Izquierdo-Villalba et al., 2021), it is still possible that SAM may offer results that is more compatible with observational data points. Nevertheless, some realizations can still yield GWB signals that reach detection levels in specific frequency bins. Furthermore, it is also possible that the GWB signal detected is a complex superposition of signals from multiple sources across the universe. While SMBBHs are known to be significant contributors to this background, they are not the sole constituents. Other physical processes with cosmological origins, such as cosmic string or cosmic phase transition, may also imprint their signatures on the GWB, thereby complicating the interpretation of the observed signals and may also contribute to the discrepancy. However, it is important to emphasize that, despite these complexities, our current projected value of the GWB amplitude falls well within the theoretical range of [$1\times 10^{-16}$, $5\times 10^{-15}$] given by existing literature in the $\tau_{H}=0.1$ and $5\,\mathrm{Gyr}$ cases.
The general physical scenario we considered in constructing light-cone SMBBH events is that we assume SMBBHs will evolve to a GW-dominated regime at 1 nHz following the mergers of their progenitor galaxies through hardening processes characterized by a fixed hardening timescale $\tau_{H}$. Since current $N$-body simulations combined with semi-analytical models cannot provide accurate information about the frequencies and orbits of SMBBH systems, we also made the simplifying assumption of circular orbital evolution from 1 nHz onward driven exclusively by GW emission. Although this is a rather simplified representation of the actual SMBBH evolution process, we’ve employed this relatively simple model to efficiently construct multiple realizations of the light-cone SMBBH dataset. In the future, we’ll consider refine our light-cone construction method by incorporating non-circular orbits and more detailed evolution modeling.
We have theoretically explored the possibility of reconstructing the large-scale structure of the universe using SMBBH events. These results may offer insights for exploring the correlations between the anisotropy of the GWB signal and the number distribution of galaxies, as is done in Semenzato et al. (2024). But it is important to recognize that since the GWB signal arises from the superposition of GW signals across different redshifts, the projection effect may weaken the correlation between GWB anisotropy and galaxy number distribution compared to the scenarios involving GW sources confined to single redshift shells considered here. In Semenzato et al. (2024), only SMBBHs within a single redshift bin were considered. It will be interesting to explore the connection between galaxy number distributions and more realistic GWB signals constructed using our complete light-cone data, as well as investigating its potential to constrain relative model parameters. We leave this to future work.